UT Mar 02 2021




Night Partition

  • Night split with LBTI - night center 00:32 - handover 00:27 to LBTO Eng LUCI-SOUL-AO
  • Moon is up effectively all night (94.6%)
  • 1st half LBTI
  • 2nd half LUCI-1 SOUL-AO
Rollover of UT Mar 01 due to failure of SX M3 moving into position

Night Plan Eng Time

AO Engineering Time - highest priority - (LBTO AO and Arcetri)
  • LUCI-SOUL AO tasks - test 6 different PAs per object
  • LUCI-SOUL Arcetri tests
    • closed loop on AO ref stars (on-axis) to collect SR and telmetry
    • one or more stars in the following categories
      • 14.5 < Rmag < 15.5 Fast Bin 3
      • 15.8 < Rmag < 16.5 Slow Bin 3
      • 16.51 < Rmag < 17.0 Fast Bin 4
      • 17.0 < Rmag < 18.0 Slow Bin 4
      • Rmag > 18.0 (as far as we can go....)
  • LUCI-SOUL ESM tasks - 1hr - only if no other full AO
    • Hoag's Object - on-axis AO Ref star (galaxy core as AO Ref star R=13.44 predicted WFSmag =14.7)
  • if cannot switch to LUCI due to M3 SX issue
    • MODS Field Correction check for IT 8315
      • 1-1.5 hours in seeing better than 1.4 arcsec, clouds don't matter much.
      • priority=medium (problem may have been with us for a year)
      • run IDL task field_collect (same as measuring field aberrations: ComTools#Field_Data_Collection)
      • do LDG and RDG in parallel (two IDL windows)
      • Use 6-8 stars (some fields only have 6 stars; start the list a second time, if you are under an hour)
      • Stone_G is well placed at RA=10:17, and other Stone / BS fields are available if earlier or later.
    • mid-z ULIRG FF0937+3857 - 2hrs MODS (Rothberg)
    • Local Mergers - 1hr per target - MODS (Rothberg)
    • LBC Solar System Objects - 1-2hrs (Veillet)
    • QSO Program - 1hr per target - MODS (Rothberg)
  • Poor Weather Backup
    • ESM data collection (Bino-LUCIs 1hr per target)
      • M67 visible from handover until 03hr
      • M92 visible from 2am through twilight
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/ESM_Testing
    • MODS and LUCI Local Mergers (Bino-MODS 90min or longer per target based on conditions) - Rothberg
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/Local_Mergers/scripts
        • MODS
          • DONE?UGC 4079 07 55 06 55 42 13 GS R=13.8
          • DONE?NGC 2744 09 04 38 18 27 37 GS R=13.9
          • NGC 6598 18 08 56 69 04 04 GS R=11.8
        • LUCI
          • DONE NGC 3310 10 38 45 53 30 05 CWL=2.27 GS R=12.5
          • NGC 4004 11 58 05 27 52 44 CWL=2.27 GS R=14.2
          • UGC 9829 15 23 01 -01 20 50 CWL=2.30 GS R=13.0
        • ICRF3 (bino-LUCIs) - Rothberg (if already switched to LUCI)
            • ESM has priority over ICRF3 - if similar conditions collect ESM data
            • /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/ICRF3
        • QSO Program - 1hr per target - MODS (Rothberg)
            • Equal priorities
            • He1029-1401
            • PG 1307+085
            • PG 1416-129
            • HB89 1635+119
      • Standard Stars (only in photometric conditions)
            • MODS primary spectrophotometric standard in all modes
              • The table of MODS primarty spectrophotometric standards is here
              • Scripts for acq and observation (all modes) are in ~modseng/modsScripts/modsSpecPhot
            • MODS imaging standards (this is better done later - the fields are only just rising at the end of the night in March)
            • LBC imaging standards
            • LUCI Standards
              • Scripts: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/LUCI_Phot_ZP
              • Best to pick one star and sample it over different airmasses
              • Otherwise, make sure to sample airmasses over delta-airmass > 0.33

Overview

NOTE: An issue with MODS2 Dichroic fault was noted tonight. Although for most data taken, the dichroic appears to have made it to position, there may have been faults not caught and MODS 2 data should be scrutinized.

Alta has a very pessimistic outlook for tonight:

And winds are expected to be high. The tree of possible plans is very complete!! And should cover all contingencies!

20210302 00:00 Waking mods and taking simSnap. Starting up LUCI's and verifying the fieldstop alignments. MODS is on obs2 (lots of x2go straggling sessions today even though I closed stragglers at the end of last night...), LUCI on robs.

Precondition violation on LUCI2 when init. Both strain gauges on. Zeroed and reinit ok

LBTI

1:17 Critical pressure loss on UBC at 20210301 15:00 today. UBC lost vacuum and is warm. Science is not possible. LBTI has assessed and determined that no engineering is possible or usable either. They will be handing the first half of the night over to LBTO engineering.

1:24 Having david move the tertiary to the LUCI/LUCI selector, then reconfiguring for MODS/MODS. Since our work is non vibration sensitive first half and LBTI is the priority, LBTI will be running their UBC pumps during the first half. Never a dull moment at the LBT. Lucikily the D plans are very thorough and flexible. We can take over tonight easily.

MODS-Bino Backup Science

1:46 Opening

2:02 Pointing and collimation. Huge temperature swing from yesterday. Ambient is -1.6 vs the -13 yesterday. Pointing and focus quite out of whack.

2:14 Preset to G191B2B. Seeing is variable ~1.4-2", closer to 2". Photometric. Doing the full spectrophotometric standard sweep here and hoping things calm a little. Dual Grating, Red Grating, blue grating, dual prism, red prism, blue prism.

Seeing variable but averaging 1.7" +/-0.4",

3:06 blew up to 4" during the dual prism, then dropped back down to around 2" for red prism on G191B2B

3:25 Seeing has been holding ~2" or just below. Presetting to V1112Per for Marks Science. Using 1.2" slit, PA97. Seeing 1.33-2.6", looks to be more variable on the left.

No errors or complaints but the first blue exposures never started on MODS2, but the seconds did. so MODS2 blue is shy one blue exposure. I don't see anythign in the MODS2 blue exposure mods2b.20210302.0015-016 I do not see the target. Looke like these were impacted by the dichroic fault (Red only dichroic rather than dual) restarting the script. Resent inst config, locked imcs and blue looked ok so that looks like it, so 0017 onward ok

Seeing ~1.8-2" on the guider 2.2-2.5" on the DIMM

4:16 Presettting off to UGC4079 for some Local Merger science. Seeing ~2". Limits reached on the right side. Must clear the right. Pointing check needed in the end so my alignment is lost.

David is struggling ot have a successful active preset. We are just going to go collimate on our field.

4:38 Resending preset to UGC4079. Seeing has improved dramatically and looks to be ~1" now. with a few brief blowups of 1.4. If it continues on this trend we will try to execute the field correction after this. Again MODS2 didn't start and it looks liek it did not go through the IMCS lock. So I restarted the mods2 dual grating script ~2 min later than mods1.

5:16 Seeing was holding at 1" but has blown up to 2" again.

Seeing has been highly variable. 0.8-2.2. Probably averaging <1.2 for the mostpart until 5:56 when I made a plan to go to one of Barry's targets that needed those conditions. Then it blew up... Seeing now 2.25" ont he guider. Going to execute another Local Merger target.

6:02 Slewing to NGC2744. During Acq, Seeing varied from 1.2-2.3"

M2> INSTCONFIG red imaging
STATUS: WARNING: **FAULT** - DICHROIC DICHROIC=FAULT Mechanism out-of-positionDONE: INSTCONFIG completed with 1 faults

M2> INSTCONFIG red imaging
STATUS: WARNING: **FAULT** - DICHROIC DICHROIC=FAULT Mechanism out-of-positionDONE: INSTCONFIG completed with 1 faults
reset dichroic

several IMCS Red lock failures on MODS2, but once i brought up the gui's too look at the laser spots succeeded first try...

There have been a number of dichroic failts tonight. I have written up IT 8333

7:34 My neighbourhood network just had an outage. Fun! Barry was online and jumped in to cover for the 10 minutes I was down. Thank you Barry!!

7:36 5" on the Dimm. 3" for the most part

MODS-2 data all night may have been affected by this dichroic issue, PIs will need to double check data

LUCI-Bino Backup Science

7:55 Reconfig LUCI/LUCI

LUCI2 rotator was not moving. David cycled and was able to get it to move again.

8:41 We can't collimate or guide on an 11th mag guide star. Seeing 5". We are going to hold for a minute or two.

8:46 Getting too late for Arp156 and going to try NGC3310. Slewing to telluric HD95126. Seeing 2.6-3.3" on the guider.

LUCI1:-9.9706, -1.3417, LUCI2 -9.4653, -0.866. LUCI2 counts seemed lower than I expected..but with 3"... Took an additional alignment at the end and all looked good so moving on. And yes...I did that...

9:10 Seeing 3.2" Slewed to NGC3310

LUCI1 0.5513, -0.5089, LUCI2 0.7547, 0.4772

Seeing is 3.7" on the guider

5.66" on the guider...thats a new record

10:17 Seeing looks like it might be improving. 1.5" on the guiders.

10:36 Preset to PG1444 Telluric HD128039 (QSO test for HBand). Central wavelength adj to 2.12

AO Testing

11:05 Seeing has improved. Assembling the AO Army and abandoning the QSO test target

UT 11:46 - J.Power lost internet (wki log may be lost between 11:05 and 11:46) - B.Rothberg jumped on.

-trying Star 25 Rmag=15.5 PA=60 Rot Angle = 312

-camera railed at PA=0 or Rot=246,

Jenny is back online.

-WFSmag = 16.9

RunAO completed - but hasn't turned it back over to LUCI

UT 11:56 - LUCI still thinks AO is trying to close the loop

-Had to click ABORT in ther Other Menu in LUCI-GUI and that got us out of the stuck preset

OBS3 has slowed down - stopped sharing on zoom, i can't

-back

UT 12::04 - 25 seconds for exposure time, ~800 counts in the peak

-taking 8x25sec exposure with telemetry.

UT 12:10 - Shell RIP'd AdSec in failure

UT 12:15 - resending with Rmag = 16.9 in it to avoid re-optimization at a different bin

UT 12:24 - E.Pinna calling it - Seeing 2.45"@Zenith on the DIMM - having trouble keeping loops closed

12:25 reauthorizing for bino luci.

LUCI-Bino Backup Science Again

12:34 Slewing to PG1444 for some last minute science. David restored IE and CA values. But pointing was still off so need to execute a pointing check.

Starting 1.25" Seeing

13:12 Only executed portion of the desired # of exposures. Getting birght so we are done. Called the mountain, we are ready to close up.

David is having trouble retracting on of the pins. Rick and david are working on the issue.

14:08 Starting some calibrations:

LUCI cals for Local Mergers data G210 2.27 K 0.5" Slit
LUCI Cals for PG1444 G210 2.12 K 0.5" Slit (incomplete). ** note that the wavelength in the OT was fixed

MODS Cals - 5arcsec dual grating flats, 5arcsec red grating flats

-- BarryRothberg - 01 Mar 2021
Topic revision: r10 - 02 Mar 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback